OpenStudy (anonymous):

Recently, the total human population reached 2051660291567 and was still growing exponentially. On a world-wide basis, the birth rate was 355 individuals per thousand people per year and the death rate was 180 individuals per thousand people per year. What was the annual rate of population growth? If you would like, you can enter your answer using scientific notation. For example, 2070000 would be entered as 2.07E06.

OpenStudy (anonymous):

annual growth rate = (annual birth rate - annual death rate) x 100/1 AGR = (355/1000 - 180/1000) x100/1 AGR = (0.355- 0.18) x 100/1 AGR = (0.175) x 100/1 AGR = 17.5%